Quiz 2 A track star runs two races on a certain day. The probability that he wins the first race is 0.7, the probability that he win the second race is 0.6, and the probability that he wins both races is 0.5 Find the probability of the following events. (5 points) (a) he wins at least one race; (5 points) (b) he wins exactly one race; (2 bonus points) (c) he wins the second race given that he wins the first one?
Answer #1

Ans -(0-3)メ(0.4) ュ0.88 (b) HHe win one saue 0(o 4)+0-3 0-6 = 0.28 + 0.18 C) = 0.3x04 0.12